# An Overview of 7726 User Reports: Uncovering SMS Scams and Scammer Strategies

### v1

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.240987 www.ndss-symposium.org
@@ I Introduction
There has been a recent surge in SMS scams worldwide [1, 2], with over 300kk fraudulent texts sent everyday [3]. Unlike phishing, where data is easily accessible through aggregators such as OpenPhish [4], Phishtank [5] and APWG eCX [6], studying SMS scams is a lot harder due to unavailability of updated public data. Blocking scam texts is not universally implemented by mobile network operators and the limited metadata available in SMS — sender ID and timestamp, makes detection challenging. In 2024, users in the US lost \$470mm to text scams [7], including \$129kk attributed to just toll-related smishing campaigns [8]. In the same period, the UK reported an 8% increase in SMS scams [9], contributing to more than £162mm in losses from telecommunications-enabled authorized push payment (APP) fraud [10]. While Australia recorded over \$14mm in losses due to smishing in 2024 [11], fraudsters stole over \$4.2mm in just three months from users in New Zealand [12]. Notably, most scams are delivered via text messages and phone calls [9].
@@
Despite the substantial financial losses caused by SMS scams, most government (or commercial) reports fail to analyze and publish detailed reports on SMS scams. For example, the FBI IC3 annual report groups smishing, vishing, and phishing into one, without differentiation [13]. The lack of insights into the SMS scam ecosystem can be attributed to the difficulties organizations face in data sharing.
@@
Government telecom regulators in the US, Canada, UK, New Zealand, and Australia work with mobile network operators to combat the increasing amount of telecom-related scams. To this end, mobile network operators have implemented SMS firewall filters and detection systems [14, 15, 16, 17] to stop SMS scams, including smishing (SMS phishing). Similar to phishing, this has become a cat-and-mouse game where scammers create new URLs [18], and mobile network operators, along with threat intelligence organizations, detect and block known URLs and text messages containing them. Despite the mobile network operator’s filters blocking malicious messages, scammers evade detection by changing sender IDs, message text, and URLs. Prior work has investigated the texts blocked by these filters [19] and we lack insights into scam texts that evade detection and reach users.
@@
As mobile network operators continue to block scams, threat actors find new tactics to evade detection and deceive users. One such technique includes sending a text message and asking users to call or text back. Others found that prompting users to reply and click on a URL increases the odds of a response to a smishing text [20]. While 7 in 10 people receive a suspicious text [21], 1 in 10 fall victim [22]. Controlled studies show at least 17% of participants fall for a smishing attack [20, 23] and users focus on the SMS content rather than the sender ID to identify smishing [24]. To devise effective countermeasures, it is essential to understand the various successful types of scams and how scammers deceive their victims.
@@
Mobile network operators in the US, Canada, UK, and New Zealand run a special reporting service called 7726 (‘SPAM’ on a mobile keypad), which allows users to report suspicious calls and SMS messages [25, 26, 27, 28]. Reported messages have evaded the mobile network operator’s filters, and this service helps mobile network operators update their detection rules with new threats. Mobile network operators collaborate with Google and Apple to integrate one-click reporting into the 7726 data feed. As users confuse legitimate and illegitimate text messages [29], the reports submitted to 7726 are undifferentiated and contain a mix of spam, scams, and legitimate texts.
@@
Research Gap. Mobile network operators have implemented an SMS detection system to reduce the amount of smishing received by end users. While previous research has examined these blocked SMS texts to study an individual scam [30] or identify scam types [19], it is essential to understand the text messages that bypass the filters and get delivered to the users. The data collected by previous studies through online forums [31], public online SMS gateways [32] or crowdsourcing [33] does not differentiate between scams and spam. This distinction is vital to build better mechanisms to thwart monetarily damaging cybercrime, as scams pose a significantly greater threat — they manipulate trust to inflict harm, making their detection and prevention a priority.
@@ II Background
With online messaging communication channels, there has been an overall decline in the total number of text messages sent/received. However, there has been an uptick in SMS texts for essential services such as order updates, bank transactions, and medical services. As of 2024, 97% of adults in the UK own or have access to mobile telephone [34], with 89.6mm active mobile subscriptions at the end of Q2 2024 [35].
@@
Recent studies on suspicious SMS messages fail to differentiate between these two types of text messages [32, 31, 33, 36]. The messages collected in previous works consist of old URLs that cannot be resolved now, and the available datasets do not include the redirected URL. Using previously collected data to differentiate spam from scams makes it challenging; §VI further discusses related academic work.
@@
There exist six known SMS scam types – (1) Wrong Number, (2) Hi Mum/Dad, (3) Delivery, (4) Banking, (5) Telecom, and (6) Government [19]. Based on these scam types, Mobile network operators block text messages which their XDRs can identify. However, many messages (which contain unknown scams) successfully evade the XDRs’ filters. Our research focuses on the SMS text messages delivered to final users, evading the mobile network operator’s detection capabilities. We provide a methodological framework that systematically groups user reports, differentiates between spam and scam, and understands the lures scammers use to deceive victims into different scams.
@@ III-B Data Enrichment
We investigate the mobile numbers users report via Home Location Register (HLR) lookup [37]. HLR lookup validates the queried phone number and provides the country to which the phone number belongs. It further provides details about a mobile number – current mobile network operator, original mobile network operator, and its current status (dead or alive). Current mobile network operator is the name of a network where the phone number is currently assigned, and the original mobile network operator is the network assigned to this telephone number range. These can change over time due to company name changes, mergers, acquisitions, and re-allocation of a number range [38].
To identify the country, the current and original mobile network operator, and its current status, we partner with Stour Marine, who offered HLR lookups [39]. We make these lookups against all newly used phone numbers as soon as we receive the user reports. We additionally query all phone numbers that were live the previous week. We combine these results to calculate each mobile number’s lifetime and find the various mobile network operators scammers abuse.
@@
We conduct survival analysis, which helps reveal patterns of activeness of mobile numbers scammer abuse. In particular, this technique considers intermittently unavailable data points to be “right-censored”; we only monitor these mobile numbers until July 18, 2024, and censor those still active on that date. We use a Kaplan-Meier estimator [40] to estimate the survival function $S(t)$ from the lifetime we observe in the data. Intuitively, this measure illustrates the fraction of mobile numbers that become inactive after a given date. This helps us estimate the lifetime of these numbers by using the probability of a mobile number abused by a scammer being active after $x$ weeks.
URLs. The majority of user-reported SMS texts contain URLs. We query all collected URLs on VirusTotal [41], an antivirus aggregator service with over 70 antivirus scanners. We extract threat intelligence from VirusTotal to determine if URLs are malicious. We consider a URL malicious if it has been flagged as such by at least one scanner.
@@
Passive DNS: Spamhaus [42] provides us with passive DNS (pDNS) API that returns the first time Spamhaus saw a domain and the IP addresses it resolved to in the last year [43].
@@
We identify registration data for each domain using a WHOIS API service. We query WhoisXMLAPI [44] at the time users report SMSs, which allows us to obtain current registration data for each domain name.
Autonomous Systems (ASes): We query the IP addresses returned by Spamhaus against IPinfo’s database [45] to identify the corresponding Autonomous System (AS) to the IP and the IP’s geographical location [46]. We note that IPinfo is unable to identify the AS for 715 IP addresses.
@@ III-C Text Classification
First, we run all unique suspicious text messages with our prompt via OpenAI’s model to categorize the text messages into spam. Next, we query the URLs from the suspicious text messages for antivirus detection on VirusTotal and query all the identified malicious domains against the domains in the text of all unique suspicious text messages with URLs. Lastly, we classify suspicious text messages with and without URLs using OpenAI’s GPT-4o model into various scam types. In particular, we identify the six known SMS scam types in the UK introduced in Section II, i.e.: (1) Wrong Number, (2) Hi Mum/Dad, (3) Delivery, (4) Banking, (5) Telecom, and (6) Government [19]. The remaining suspicious text messages with URLs where VirusTotal does not flag the URL are ‘Unknown.’
@@
Complexity Analysis. Gunning Fog Index helps explain the educational level required to understand the scam text message the Gunning fog index [47]. The scam text messages collected in our dataset are in English, as they target users in the UK. We tokenize the text of the SMS messages reported by users and remove the stop words. Stop words are common words that do not provide meaningful information about the topic. To calculate the complex words, we use the conventional criteria of the Gunning Fog Index and filter out the words with less than three syllables. We additionally remove proper nouns, such as brand names being impersonated. The formula to calculate the Gunning Fog Index is as follows:
@@
Scam Lure Analysis. We want to understand the various lures that scammers use to deceive victims into taking actions mentioned in a scam message. Towards this end, we adopt the ontology of lures from Stajano and Wilson [48]. The n-gram analysis for multiple scam types with thousands of messages is not feasible to do manually. Thus, we use OpenAI’s GPT-4o to categorize our messages into the different lures using the prompt listed in Appendix -B.
@@ III-D Evaluation
To evaluate the performance of our methodology, we extract 384 random texts from our dataset, ensuring (95% confidence) that the random sample is representative of the complete dataset. We manually label them as an OTP, spam, or scam. Next, we extract another random sample of 384 scam texts and manually classify them into seven scam types (including other) along with the scammer lures used in each scam text. We use human-labeled texts as ground truth and calculate the inter-rater reliability (IRR) between ground truth and OpenAI’s annotation. We use Cohen’s $\kappa$ [49], a standard metric for IRR. There is near-perfect agreement for OTP and scam types and substantial agreement for spam and lures (see Table II for Cohen’s $\kappa$ coefficient and F1-score).
@@ IV-A User Reports Over Time
Takeaway. We see that users mostly submit reports between 10:00 and 21:00 BST daily. Prior work found that scammers interact with Hi mum/dad scam victims during 10:00-15:00 UK time [30], indicating that users report SMS messages or calls without significant delays.
@@ IV-B1 Suspicious Text Message Classification
These messages are generally unsolicited marketing messages sent by companies or unknown senders. Spam is often annoying and reduces the availability of messages on a user’s mobile phone, which can prompt users to report them to 7726. Some of these reports might be erroneous. As reported by Ofcom, the built-in reporting function on users’ mobile phones was the most used channel to report suspicious messages [50]. The user interface has ‘delete’ and ‘delete and report’ buttons next to each other, which could confuse users and send their reports to 7726 instead of simply deleting the text message.
@@
Unknowns. We label the remaining 123,428 (23.3%) unique suspicious text messages with URLs as ‘Unknown’. Considering that these messages have been reported by users and flagged as suspicious, we see them as potentially malicious rather than benign due to the various evasion techniques deployed by phishing websites [51] and the low recall of AV vendors in detecting phishing websites [52, 36].
@@ IV-B2 Scam Type Classification
We classify scam text messages into six known SMS scam types [19] and identify six new scam types using our methodological process described in §III-C. Table V outlines messages by category. Out of the 119,398 scam texts overall, we discover that the most popular scam type is the Wrong Number scam (16.36%), followed by Banking (9.14%) and Delivery/Parcel (6.81%) scams. We plot the distribution of the these scams over time in Fig. 3(a).
@@
Known [19] Wrong Number 34,863 86 34,949 16.36
@@
From the new scam categories, we find that job-related scams (5.12%) are the most reported category, followed by debt (4.85%) and appointment-related scams (2.13%). Job-related scams lure victims by providing fake employment opportunities [53]. Debt scams are text messages that ask users to pay an outstanding debt balance. Appointment scams are similar to appointment reminders, asking users to call or text back on the phone number but then luring victims into transferring funds or stealing financial or personal details [54, 55]. Example of job-related scams is (cf. Appendix -C for other types):
@@
Some scam texts contain a malicious URL and are known as smishing or SMS phishing. These messages contain a malicious URL enticing a user to click; this redirects to a phishing page impersonating a brand which deceives victims into providing personal or financial details. Unsurprisingly, Delivery/Parcel (12,725) and Banking (4,784) impersonation scams are the top two categories for URL-based scams (Table V). From the new categories, for URL-based scams, we find Debt (3,185) and Utility (1,355) scams to be the top two. Utility scams impersonate a utility company, such as gas or water, and aim to steal users’ details and banking credentials [56]. Here is an example Banking scam text with a URL where scammers try to lure victims into providing credentials:
@@
We do not expect conversational scams to contain URLs. However, we see a handful of these (86 Wrong Number, 190 Hi Mum/Dad) with URLs, overwhelmingly from scammers attempting to move the conversation to online messaging platforms, e.g., WhatsApp. Contrary to URL-based scams, scams without URLs feature Wrong Number scams (34,863), followed by Banking scams. These are more difficult to identify than a malicious URL in a smishing text. While Hi Mum/Dad scams are a well-known authorized push payment (APP) fraud which cause significant financial loss to victims [30], these are actively blocked by mobile network operators in the UK [19] and thus lower in ranking by volume (6,604). Investigating banking scam messages without a URL, we find that these messages mention an OTP or a transaction and ask the victim to call the provided phone number. For example,
@@
This indicates a rise in text and call-back scams. Here, scammers lure victims through text messages by asking them to text or call the phone number provided in the SMS text. These messages often include a OTP or mention a fake transaction, instructing recipients to call or text a provided phone number if they did not request the OTP or initiate a transaction. We do not directly engage with phone numbers provided, either by texting or calling; telephony honeypots present a potential avenue for deeper investigation [57] discussed in §V.
We note new scam categories have higher volumes for non-URL-based scams, particularly Job scams (10,407) and Debt-based scams (7,188) [58, 59, 53]. Job scams exploit users’ vulnerability by contacting them about unrealistic job opportunities, offering high salaries to lure them into a scam [60, 61]. We plot the distribution of the new scam types over time in Fig. 3(b).

### v2

@@ I Introduction
There has been a recent surge in SMS scams worldwide [trendmicro_scam, techradar_smsscam], with over 300kk fraudulent texts sent everyday [proofpoint_mwc]. Unlike phishing, where data is easily accessible through aggregators such as OpenPhish [openphish], Phishtank [phishtank] and APWG eCX [apwg_ecx], studying SMS scams is a lot harder due to unavailability of updated public data. Blocking scam texts is not universally implemented by mobile network operators and the limited metadata available in SMS — sender ID and timestamp, makes detection challenging. In 2024, users in the US lost \$470mm to text scams [ftc_smsscams], including \$129kk attributed to just toll-related smishing campaigns [fbi_ic3:24]. In the same period, the UK reported an 8% increase in SMS scams [gasa_uk], contributing to more than £162mm in losses from telecommunications-enabled authorized push payment (APP) fraud [ukfinance_2025]. While Australia recorded over \$14mm in losses due to smishing in 2024 [scamwatch_aus], fraudsters stole over \$4.2mm in just three months from users in New Zealand [nz_loss]. Notably, most scams are delivered via text messages and phone calls [gasa_uk].
@@
Despite the substantial financial losses caused by SMS scams, most government (or commercial) reports fail to analyze and publish detailed reports on SMS scams. For example, the FBI IC3 annual report groups smishing, vishing, and phishing into one, without differentiation [fbi_ic3]. The lack of insights into the SMS scam ecosystem can be attributed to the difficulties organizations face in data sharing.
@@
Government telecom regulators in the US, Canada, UK, New Zealand, and Australia work with mobile network operators to combat the increasing amount of telecom-related scams. To this end, mobile network operators have implemented SMS firewall filters and detection systems [canada:mno_filter, spamshield_mavenir, nz:mno_filter, aus:mno_filter] to stop SMS scams, including smishing (SMS phishing). Similar to phishing, this has become a cat-and-mouse game where scammers create new URLs [agarwal25:weis], and mobile network operators, along with threat intelligence organizations, detect and block known URLs and text messages containing them. Despite the mobile network operator’s filters blocking malicious messages, scammers evade detection by changing sender IDs, message text, and URLs. Prior work has investigated the texts blocked by these filters [agarwal_imc24] and we lack insights into scam texts that evade detection and reach users.
@@
As mobile network operators continue to block scams, threat actors find new tactics to evade detection and deceive users. One such technique includes sending a text message and asking users to call or text back. Others found that prompting users to reply and click on a URL increases the odds of a response to a smishing text [lutforsmishing_interview]. While 7 in 10 people receive a suspicious text [ofcom_21], 1 in 10 fall victim [cifas]. Controlled studies show at least 17% of participants fall for a smishing attack [lutforsmishing_interview, timko2025understanding] and users focus on the SMS content rather than the sender ID to identify smishing [smishinguserstudy_soups24]. To devise effective countermeasures, it is essential to understand the various successful types of scams and how scammers deceive their victims.
@@
Mobile network operators in the US, Canada, UK, and New Zealand run a special reporting service called 7726 (‘SPAM’ on a mobile keypad), which allows users to report suspicious calls and SMS messages [ofcom_7726, att_7726, nz_7726, ca_7726]. Reported messages have evaded the mobile network operator’s filters, and this service helps mobile network operators update their detection rules with new threats. Mobile network operators collaborate with Google and Apple to integrate one-click reporting into the 7726 data feed. As users confuse legitimate and illegitimate text messages [timko2023more], the reports submitted to 7726 are undifferentiated and contain a mix of spam, scams, and legitimate texts.
@@
Research Gap. Mobile network operators have implemented an SMS detection system to reduce the amount of smishing received by end users. While previous research has examined these blocked SMS texts to study an individual scam [agarwal_usenix25] or identify scam types [agarwal_imc24], it is essential to understand the text messages that bypass the filters and get delivered to the users. The data collected by previous studies through online forums [tang:ccs22], public online SMS gateways [nahapetyan2024sms] or crowdsourcing [smishtank_timko24] does not differentiate between scams and spam. This distinction is vital to build better mechanisms to thwart monetarily damaging cybercrime, as scams pose a significantly greater threat — they manipulate trust to inflict harm, making their detection and prevention a priority.
@@ II Background
With online messaging communication channels, there has been an overall decline in the total number of text messages sent/received. However, there has been an uptick in SMS texts for essential services such as order updates, bank transactions, and medical services. As of 2024, 97% of adults in the UK own or have access to mobile telephone [ofcom_mobileaccess], with 89.6mm active mobile subscriptions at the end of Q2 2024 [ofcom_activesubscription].
@@
Recent studies on suspicious SMS messages fail to differentiate between these two types of text messages [nahapetyan2024sms, tang:ccs22, smishtank_timko24, srinivasan2016understanding]. The messages collected in previous works consist of old URLs that cannot be resolved now, and the available datasets do not include the redirected URL. Using previously collected data to differentiate spam from scams makes it challenging; §VI further discusses related academic work.
@@
There exist six known SMS scam types – (1) Wrong Number, (2) Hi Mum/Dad, (3) Delivery, (4) Banking, (5) Telecom, and (6) Government [agarwal_imc24]. Based on these scam types, Mobile network operators block text messages which their XDRs can identify. However, many messages (which contain unknown scams) successfully evade the XDRs’ filters. Our research focuses on the SMS text messages delivered to final users, evading the mobile network operator’s detection capabilities. We provide a methodological framework that systematically groups user reports, differentiates between spam and scam, and understands the lures scammers use to deceive victims into different scams.
@@ III-B Data Enrichment
We investigate the mobile numbers users report via Home Location Register (HLR) lookup [hlrlookup]. HLR lookup validates the queried phone number and provides the country to which the phone number belongs. It further provides details about a mobile number – current mobile network operator, original mobile network operator, and its current status (dead or alive). Current mobile network operator is the name of a network where the phone number is currently assigned, and the original mobile network operator is the network assigned to this telephone number range. These can change over time due to company name changes, mergers, acquisitions, and re-allocation of a number range [hlrlookup_api].
To identify the country, the current and original mobile network operator, and its current status, we partner with Stour Marine, who offered HLR lookups [stourmarine]. We make these lookups against all newly used phone numbers as soon as we receive the user reports. We additionally query all phone numbers that were live the previous week. We combine these results to calculate each mobile number’s lifetime and find the various mobile network operators scammers abuse.
@@
We conduct survival analysis, which helps reveal patterns of activeness of mobile numbers scammer abuse. In particular, this technique considers intermittently unavailable data points to be “right-censored”; we only monitor these mobile numbers until July 18, 2024, and censor those still active on that date. We use a Kaplan-Meier estimator [kaplan1958nonparametric] to estimate the survival function $S(t)$ from the lifetime we observe in the data. Intuitively, this measure illustrates the fraction of mobile numbers that become inactive after a given date. This helps us estimate the lifetime of these numbers by using the probability of a mobile number abused by a scammer being active after $x$ weeks.
URLs. The majority of user-reported SMS texts contain URLs. We query all collected URLs on VirusTotal [virustotal], an antivirus aggregator service with over 70 antivirus scanners. We extract threat intelligence from VirusTotal to determine if URLs are malicious. We consider a URL malicious if it has been flagged as such by at least one scanner.
@@
Passive DNS: Spamhaus [spamhaus] provides us with passive DNS (pDNS) API that returns the first time Spamhaus saw a domain and the IP addresses it resolved to in the last year [spamhaus_pdns].
@@
We identify registration data for each domain using a WHOIS API service. We query WhoisXMLAPI [whoisxml] at the time users report SMSs, which allows us to obtain current registration data for each domain name.
Autonomous Systems (ASes): We query the IP addresses returned by Spamhaus against IPinfo’s database [ipinfo] to identify the corresponding Autonomous System (AS) to the IP and the IP’s geographical location [ipinfo_db]. We note that IPinfo is unable to identify the AS for 715 IP addresses.
@@ III-C Text Classification
First, we run all unique suspicious text messages with our prompt via OpenAI’s model to categorize the text messages into spam. Next, we query the URLs from the suspicious text messages for antivirus detection on VirusTotal and query all the identified malicious domains against the domains in the text of all unique suspicious text messages with URLs. Lastly, we classify suspicious text messages with and without URLs using OpenAI’s GPT-4o model into various scam types. In particular, we identify the six known SMS scam types in the UK introduced in Section II, i.e.: (1) Wrong Number, (2) Hi Mum/Dad, (3) Delivery, (4) Banking, (5) Telecom, and (6) Government [agarwal_imc24]. The remaining suspicious text messages with URLs where VirusTotal does not flag the URL are ‘Unknown.’
@@
Complexity Analysis. Gunning Fog Index helps explain the educational level required to understand the scam text message the Gunning fog index [gunning1969fog]. The scam text messages collected in our dataset are in English, as they target users in the UK. We tokenize the text of the SMS messages reported by users and remove the stop words. Stop words are common words that do not provide meaningful information about the topic. To calculate the complex words, we use the conventional criteria of the Gunning Fog Index and filter out the words with less than three syllables. We additionally remove proper nouns, such as brand names being impersonated. The formula to calculate the Gunning Fog Index is as follows:
@@
Scam Lure Analysis. We want to understand the various lures that scammers use to deceive victims into taking actions mentioned in a scam message. Towards this end, we adopt the ontology of lures from Stajano and Wilson [StajanoFrank2011_principles]. The n-gram analysis for multiple scam types with thousands of messages is not feasible to do manually. Thus, we use OpenAI’s GPT-4o to categorize our messages into the different lures using the prompt listed in Appendix -B.
@@ III-D Evaluation
To evaluate the performance of our methodology, we extract 384 random texts from our dataset, ensuring (95% confidence) that the random sample is representative of the complete dataset. We manually label them as an OTP, spam, or scam. Next, we extract another random sample of 384 scam texts and manually classify them into seven scam types (including other) along with the scammer lures used in each scam text. We use human-labeled texts as ground truth and calculate the inter-rater reliability (IRR) between ground truth and OpenAI’s annotation. We use Cohen’s $\kappa$ [cohen1960coefficient], a standard metric for IRR. There is near-perfect agreement for OTP and scam types and substantial agreement for spam and lures (see Table II for Cohen’s $\kappa$ coefficient and F1-score).
@@ IV-A User Reports Over Time
Takeaway. We see that users mostly submit reports between 10:00 and 21:00 BST daily. Prior work found that scammers interact with Hi mum/dad scam victims during 10:00-15:00 UK time [agarwal_usenix25], indicating that users report SMS messages or calls without significant delays.
@@ IV-B1 Suspicious Text Message Classification
These messages are generally unsolicited marketing messages sent by companies or unknown senders. Spam is often annoying and reduces the availability of messages on a user’s mobile phone, which can prompt users to report them to 7726. Some of these reports might be erroneous. As reported by Ofcom, the built-in reporting function on users’ mobile phones was the most used channel to report suspicious messages [ofcom_userexperience]. The user interface has ‘delete’ and ‘delete and report’ buttons next to each other, which could confuse users and send their reports to 7726 instead of simply deleting the text message.
@@
Unknowns. We label the remaining 123,428 (23.3%) unique suspicious text messages with URLs as ‘Unknown’. Considering that these messages have been reported by users and flagged as suspicious, we see them as potentially malicious rather than benign due to the various evasion techniques deployed by phishing websites [doupe_usenix20] and the low recall of AV vendors in detecting phishing websites [peng_imc19, srinivasan2016understanding].
@@ IV-B2 Scam Type Classification
We classify scam text messages into six known SMS scam types [agarwal_imc24] and identify six new scam types using our methodological process described in §III-C. Table V outlines messages by category. Out of the 119,398 scam texts overall, we discover that the most popular scam type is the Wrong Number scam (16.36%), followed by Banking (9.14%) and Delivery/Parcel (6.81%) scams. We plot the distribution of the these scams over time in Fig. 3(a).
@@
Known [agarwal_imc24] Wrong Number 34,863 86 34,949 16.36
@@
From the new scam categories, we find that job-related scams (5.12%) are the most reported category, followed by debt (4.85%) and appointment-related scams (2.13%). Job-related scams lure victims by providing fake employment opportunities [which_scams]. Debt scams are text messages that ask users to pay an outstanding debt balance. Appointment scams are similar to appointment reminders, asking users to call or text back on the phone number but then luring victims into transferring funds or stealing financial or personal details [appointment_scam, fakeappointment_scam]. Example of job-related scams is (cf. Appendix -C for other types):
@@
Some scam texts contain a malicious URL and are known as smishing or SMS phishing. These messages contain a malicious URL enticing a user to click; this redirects to a phishing page impersonating a brand which deceives victims into providing personal or financial details. Unsurprisingly, Delivery/Parcel (12,725) and Banking (4,784) impersonation scams are the top two categories for URL-based scams (Table V). From the new categories, for URL-based scams, we find Debt (3,185) and Utility (1,355) scams to be the top two. Utility scams impersonate a utility company, such as gas or water, and aim to steal users’ details and banking credentials [utility_scam]. Here is an example Banking scam text with a URL where scammers try to lure victims into providing credentials:
@@
We do not expect conversational scams to contain URLs. However, we see a handful of these (86 Wrong Number, 190 Hi Mum/Dad) with URLs, overwhelmingly from scammers attempting to move the conversation to online messaging platforms, e.g., WhatsApp. Contrary to URL-based scams, scams without URLs feature Wrong Number scams (34,863), followed by Banking scams. These are more difficult to identify than a malicious URL in a smishing text. While Hi Mum/Dad scams are a well-known authorized push payment (APP) fraud which cause significant financial loss to victims [agarwal_usenix25], these are actively blocked by mobile network operators in the UK [agarwal_imc24] and thus lower in ranking by volume (6,604). Investigating banking scam messages without a URL, we find that these messages mention an OTP or a transaction and ask the victim to call the provided phone number. For example,
@@
This indicates a rise in text and call-back scams. Here, scammers lure victims through text messages by asking them to text or call the phone number provided in the SMS text. These messages often include a OTP or mention a fake transaction, instructing recipients to call or text a provided phone number if they did not request the OTP or initiate a transaction. We do not directly engage with phone numbers provided, either by texting or calling; telephony honeypots present a potential avenue for deeper investigation [miramirkhani2016_techscam] discussed in §V.
We note new scam categories have higher volumes for non-URL-based scams, particularly Job scams (10,407) and Debt-based scams (7,188) [debt_scamwithouturl, job_scam, which_scams]. Job scams exploit users’ vulnerability by contacting them about unrealistic job opportunities, offering high salaries to lure them into a scam [job_scamwhich, ftc_jobscams]. We plot the distribution of the new scam types over time in Fig. 3(b).